# Verifiable mission planning for space operations

### v2

@@ 4.1 GRACE-FO mission
The NASA Gravity Recovery and Climate Experiment (GRACE) mission, launched in 2002, was a NASA mission that sought to obtain high-accuracy estimates of Earth’s gravity field via high-resolution spatial and temporal measurements [16]. The mission consisted of a pair of spacecraft, named GRACE-1 and GRACE-2, operating in near-circular orbits at an inclination of $89.5$ degrees and an altitude of approximately $500$ kilometers.
The GRACE-FO (Gravity Recovery and Climate Experiment Follow-On) mission, launched in May 2018, continues the legacy of the original GRACE mission to measure Earth’s gravity field with unprecedented precision. By monitoring variations in the inter-satellite range using K-Band Ranging and Laser Ranging Interferometer technologies, GRACE-FO provides critical data on global surface mass changes.
One major challenge arises from the failure of the accelerometer on GRACE-FO 2 (GF2), rendering it unable to directly measure non-gravitational forces. To compensate, data from the accelerometer of GRACE-FO 1 (GF1) is transplanted to GF2, relying on the assumption that the environmental conditions experienced by both satellites are nearly identical within a 24-30-second delay. Although this method captures over $99\%$ of the non-gravitational forces, it introduces signal-proportional errors amplified in high-drag environments, degrading the accuracy of derived gravity fields. Furthermore, the exact nature of the current solar cycle is uncertain; it could exhibit high, moderate, or low activity, each scenario presenting different levels of atmospheric drag. To address these uncertainties, we aim to develop a robust strategy that accounts for the different scenarios of solar activity and their influence on atmospheric drag, ensuring reliable mission operations and data quality.
Another critical issue involves a small leak in the cold-gas thrusters used for orbit maintenance. Every use of the thrusters risks worsening the leak, potentially increasing the rate of fuel loss and compounding the challenge of preserving enough fuel for critical operations. This creates a dynamic trade-off between performing necessary station-keeping maneuvers and avoiding actions that might accelerate the leak, further threatening the mission’s longevity and scientific output.
The objective is to develop a strategy to optimize GRACE-FO’s mission time and scientific output under these constraints. We explicitly account for uncertainties in space weather and fuel availability, including the impact of solar cycle variability and the thruster leak.
@@ 4.2 State Modeling
We discretize the environment of the spacecraft in a pair of dimensions. The first of these dimensions is the altitude of the spacecraft. Recall that we are primarily interested in the effect of solar flux on atmospheric drag, and that atmospheric drag only affects the semimajor axis length and the eccentricity of the spacecraft orbit. Furthermore, since the GRACE mission is nominally in a circular orbit (i.e. $e=0$), its eccentricity is unaffected by atmospheric drag and remains constant over time. Thus, we focus solely on discretizing the altitude of the spacecraft (which is a function of the semimajor axis length as given by (8)).
To construct this discretization, let $a_{min}$, $a_{max}\in\mathbb{R}\geq 0$, $a_{max}>a_{min}$, be a minimum and maximum altitude, respectively, and $N_{a}\in\mathbb{Z}$ denote the number of discrete altitudes to consider within these bounds. Denote the set of these discrete altitudes by $\mathcal{S}_{a}\triangleq\{a_{min}+\frac{R_{a}}{2},a_{min}+\frac{3R_{a}}{2},\ldots,a_{max}-\frac{R_{a}}{2}\}$, where $R_{a}\triangleq\frac{a_{max}-a_{min}}{N_{A}}$. Each altitude $s\in\mathcal{S}_{a}$ corresponds to a discrete band centered around $s$, i.e., the band $[s-\frac{R_{a}}{2},s+\frac{R_{a}}{2}]$, as shown in Figure 3. Although the true semimajor axis length of the spacecraft will not generally reside on one of the sampled altitudes $s\in\mathcal{S}_{a}$, we can instead bin it into a corresponding band (assuming that it lies within the minimum and maximum altitudes).
@@
To construct the fuel discretization, we use a similar method as for the altitude. Let $f_{min},f_{max}\in\mathbb{R}_{\geq 0}$, with $f_{max}>f_{min}$, be the minimum and maximum fuel levels, respectively, and let $N_{f}\in\mathbb{Z}$ denote the number of discrete fuel levels to consider within these bounds. We denote the set of these discrete fuel levels by $\mathcal{F}\triangleq\{f_{min}+\frac{R_{f}}{2},\;f_{min}+\frac{3R_{f}}{2},\;\ldots,\;f_{max}-\frac{R_{f}}{2}\},$ where $R_{f}\triangleq\frac{f_{max}-f_{min}}{N_{f}}.$ Following the same procedure as stated for the altitude, the true fuel level of the spacecraft is binned into the corresponding fuel band.
@@ 4.3 Reward Modeling
At a high level, the objective of the GRACE mission is to provide high-quality observations of changes in Earth’s gravitational field [17]. The quality of these observations generally increases with lower atmospheric density, which occurs at higher altitudes. Due to this fact, the reward function $R$ should be independent of the choice of the action selection and only a function of the current state (i.e., altitude) and time step of the mission. The dependence of the reward function $R$ on time could be due to, e.g., time-varying atmospheric conditions that affect the observation quality. Thus, we introduce $R(s,:,h)\triangleq R(s,a,h)$ for all $a\in\mathcal{A}$ to denote the reward (a function of the observation quality) for being in a state $s\in\mathcal{S}$ corresponding to a fixed, discrete altitude. In other words, $R(s,:,h)$ is independent of the action taken. Note that $R(s,:,h)$ should be non-negative and monotonically increasing as the corresponding altitude of $s$ increases.
An additional hindrance to observation quality is the existence of resonant orbits, which are orbits with a periodic ground track [18]. Due to the periodicity of this ground track, less of Earth’s surface is observed, reducing the quality of the global gravitational field estimation. Thus, for any discrete altitude state $s\in\mathcal{S}_{a}$ for which a resonant orbit exists in its corresponding altitude band $[s-\frac{R_{a}}{2},s+\frac{R_{a}}{2}]$, we can set $R(s,:,h)=0$. Alternatively, given periodicity of the resonant orbit $i$, $res_{i}$ (i.e., the rate at which the ground track repeats), one can scale the reward function by a multiplicative factor $\alpha<1$ to account for variations in the reduction of observation quality. The reward function can then be defined as
@@
where $\omega_{e}$ is the Earth’s rotation rate and $\dot{\Omega}$ is the nodal precession rate.
Substituting these into the resonance condition yields
@@ 4.4 Action and Fuel Cost Modeling
[…]
Associated with each action is a fuel cost, while the decision-maker must ensure that the cumulative fuel cost throughout the mission does not exceed its fuel budget. Let us denote the fuel cost when taking action $a\in\mathcal{A}$ while in a state $s\in\mathcal{S}$ by $C(s,a)$ and the fuel budget by $C_{\mathrm{max}}$. Trivially, for the action corresponding to no burn, the fuel cost is zero. For the remaining actions, there exist several options for modeling the fuel costs and the fuel budget. These options are as follows.
@@
    Maximum Altitude Change. As previously discussed, each action is associated with an increase in the operational altitude of the GRACE mission. The cost of an action is its corresponding altitude increase, i.e., $C(s,{}^{+}$$nR_{a})={}^{+}$$nR_{a}$ for all $n\in\{1,\ldots,N_{A}\}$. Similarly, the fuel budget is a maximum cumulative altitude increase over the mission horizon.
@@ 4.5 Transition Modeling
$\mathrm{Pr}(\mathrm{alt}_{h+1}\in[R_{s^{\prime}}-\frac{R_{a}}{2},R_{s^{\prime}}+\frac{R_{a}}{2}]|\mathrm{alt}_{h}=R_{s}+{}^{+}R_{a})=$ (12b)
@@
where $R_{s}$ is the altitude corresponding to state $s\in\mathcal{S}$ and ${}^{+}R_{a}$ is the altitude gain corresponding to selecting action $a\in\mathcal{A}$. Note that (12c) is obtained from (12b) by explicitly including the conditioning of the transition probability on the solar flux. Furthermore, since the state dynamics themselves are assumed to have no stochastic disturbances, $\mathrm{Pr}(\mathrm{alt}_{t+1}\in[R_{s^{\prime}}-\frac{R_{a}}{2},R_{s^{\prime}}+\frac{R_{a}}{2}]|\mathrm{alt}_{t}=R_{s}+{}^{+}R_{a},f)$ takes either a value of 00 or $1$ (i.e., it is binary). In practice, the integral over $f$ can be numerically approximated and the state transition matrix used to efficiently compute the subsequent altitude. A visualization of an example trajectory (with respect to the mission altitude) is shown in Figure 6.
@@
Note that decision are only made at the discrete time instances represented by the densely dotted vertical lines. Furthermore, the mission altitude will not generally reside on one of the discretely sampled states $s\in\mathcal{S}$. However, given a set $\mathcal{S}$ sufficiently densely sampled from the set of feasible altitudes $[a_{min},a_{max}]$, binning the actual altitude into the corresponding band should provide a good approximation of the continuous dynamics.
@@
    Collision avoidance. When the risk of collision with another spacecraft or orbital debris increases, GRACE cannot execute effective avoidance maneuvers. This risk is quantified by a collision rate, $\lambda_{c}$, whose integration over discrete time intervals yields the probability that a collision avoidance maneuver is required.
@@ 4.6 Problem Formulation
We now formalize the mission planning problem as an optimization task over a finite planning horizon. The objective is to maximize the total expected reward while ensuring that a critical safety specification is satisfied with high probability. We define a safety requirement expressed as a time logic formula that captures two key constraints:
@@
[…]
The informal safety specification can be read as follows:
”The spacecraft does not perform a maneuver more often than every 3 months, and it never drops below 300 km.”
@@ 5.2 Planning Under Uncertainties
Next, we incorporate uncertainties into the model by treating both the solar flux and the outcomes of the orbit-raising maneuvers as random variables. Figure 10 shows the evolution of the solar flux over time, with the 75% quartile range highlighted in orange; these discretized distributions are integrated into the model. We then evaluate state transitions by considering best-case, worst-case, and average scenarios based on variations in solar flux and orbit-raising cost. Figure 10 displays a range of possible trajectories for altitude and fuel consumption. As expected, the worst-case scenario leads to fuel depletion early in 2025, while the best-case scenario extends fuel availability until 2029. The final altitude varies widely—from below 300 km up to 475 km—with an average of 410 km. Although the average behavior appears to satisfy the temporal specification, this plot alone does not provide full probabilistic safety guarantees. It suggests possible violations of the requirement that the altitude remains above 300 km. We therefore now turn to safety verification.
@@
Figure 9: Altitude and fuel evolution when solving the MDP with uncertain transitions,
@@
Figure 10: Probabilistic verification for the final altitude following the policy $\pi$ using the initial condition of GRACE-FO at its launch.
@@ 5.3 Safety Verification
These results underscore the strength of the MDP framework in enforcing strict safety constraints while optimizing performance. Unlike traditional planning methods that do not explicitly verify safety through continuous feedback, the value function in this approach provides a clear measure of safety compliance. The near-zero risk of dropping below the minimum altitude and the high probability of maintaining an altitude above 425 km indicate that the adaptive policy not only meets but exceeds safety requirements.
@@
Figure 11: Altitude and fuel evolution when solving the MDP with uncertain transitions using GRACE-FO state in March 2025 as the initial condition.
@@ 5.4 Current-State Planning
The MDP formulation updates the optimal action as more accurate solar flux data becomes available. This capability enables the computation of an optimal policy for orbit-raising maneuvers that adapts to the current state of the GRACE-FO spacecraft. For example, in February 2025, when only 1.5 kg of fuel remains, the same policy computes an updated sequence of maneuvers. As shown in Figure 11, the resulting strategy clearly meets the safety specification: the policy keeps the spacecraft’s altitude well above 300 km and prevents an additional maneuver if one has been executed within the past three months, following the same principles as the full mission planning.
These current-state experiments highlight the method’s ability to adapt under critical fuel constraints. Unlike traditional approaches that schedule maneuvers at fixed intervals, the MDP-based policy adjusts its actions dynamically based on real-time state information. This adaptive behavior ensures that even when fuel is scarce, the spacecraft operates safely and efficiently. Overall, the experiments provide strong evidence that the planning method remains robust under uncertainties in the environment, effectively handling unpredictable variations in solar flux and maneuver costs while maintaining operational safety.

### v3

@@ 4.1 GRACE-FO mission
The Gravity Recovery and Climate Experiment (GRACE), launched in 2002, was a NASA mission that sought to obtain high-accuracy estimates of Earth’s gravity field from twin satellites in near-polar, near-circular orbits [15]. The GRACE-FO (Gravity Recovery and Climate Experiment Follow-On) mission, launched in May 2018, continues the legacy of the original GRACE mission to measure Earth’s gravity field with unprecedented precision. By monitoring variations in the inter-satellite range using K-Band Ranging and Laser Ranging Interferometer technologies, GRACE-FO provides critical data on global surface mass changes.
One major challenge arises from the failure of the accelerometer on GRACE-FO 2 (GF2), rendering it unable to measure non-gravitational forces directly. To compensate, the data from the accelerometer of GRACE-FO 1 (GF1) are transplanted to GF2, relying on the assumption that both satellites experience nearly identical environmental conditions within the $\sim 24–30$ s separation delay. The transplant method recovers differential non-gravitational forces with high fidelity[16] under nominal conditions, but it introduces signal-proportional errors that can reach a few percent of the drag signal in high-drag regimes, which degrades the resulting gravity fields.
Propulsive capability is another constraint. Altitude-raising was not a design objective, so the cold-gas system and fuel budget are limited for orbit-raising. In addition, thrust realization during burns is uncertain (mismatch between commanded and actual impulse), which introduces variance in achieved altitude change and propellant usage. With finite fuel, this creates a trade-off between performing raises to preserve data quality and conserving propellant for mission lifetime.
Finally, solar-cycle variability drives large, stochastic changes in upper-atmospheric density and hence drag. Periods of elevated activity accelerate orbital decay and amplify transplant-related errors; quieter periods relax both.
We seek a planning strategy that addresses three coupled challenges: the trade-off between altitude and measurement fidelity, limited fuel with uncertain thrust performance, and stochastic drag driven by solar activity. To this end, we present a finite-horizon MDP formulation with altitude and fuel as state variables, orbit-raising maneuvers as discrete actions, and transition dynamics derived from drag and thrust-realization models. The resulting formulation enables optimization of science return while ensuring compliance with safety constraints.
@@ 4.2 State Modeling
We discretize the environment of the spacecraft in a pair of dimensions. The first of these dimensions is the altitude of the spacecraft. Recall that we are primarily interested in the effect of solar flux on atmospheric drag, and that atmospheric drag only affects the semimajor axis length and the eccentricity of the spacecraft orbit. Furthermore, since the GRACE-FO mission is nominally in a circular orbit (i.e. $e=0$), its eccentricity is unaffected by atmospheric drag and remains constant over time. Thus, we focus solely on discretizing the altitude of the spacecraft (which is a function of the semimajor axis length as given by (8)).
To construct this discretization, let $a_{min}$, $a_{max}\in\mathbb{R}\geq 0$, $a_{max}>a_{min}$, be a minimum and maximum altitude, respectively, and $N_{a}\in\mathbb{Z}$ denote the number of discrete altitudes to consider within these bounds. Denote the set of these discrete altitudes by $\mathcal{S}_{a}\triangleq\{a_{min}+\frac{R_{a}}{2},a_{min}+\frac{3R_{a}}{2},\ldots,a_{max}-\frac{R_{a}}{2}\}$, where $R_{a}\triangleq\frac{a_{max}-a_{min}}{N_{A}}$. Each altitude $s\in\mathcal{S}_{a}$ corresponds to a discrete band centered around $s$, i.e., the band $s-\frac{R_{a}}{2},s+\frac{R_{a}}{2}]$, as shown in Figure 3. Although the true semimajor axis length of the spacecraft will not generally reside on one of the sampled altitudes $s\in\mathcal{S}_{a}$, we can instead bin it into a corresponding band (assuming that it lies within the minimum and maximum altitudes).
@@
To construct the fuel discretization, we use a similar method to that for the altitude. Let $f_{min},f_{max}\in\mathbb{R}_{\geq 0}$, with $f_{max}>f_{min}$, be the minimum and maximum fuel levels, respectively, and let $N_{f}\in\mathbb{Z}$ denote the number of discrete fuel levels to consider within these bounds. We denote the set of these discrete fuel levels by $\mathcal{F}\triangleq\{f_{min}+\frac{R_{f}}{2},\;f_{min}+\frac{3R_{f}}{2},\;\ldots,\;f_{max}-\frac{R_{f}}{2}\},$ where $R_{f}\triangleq\frac{f_{max}-f_{min}}{N_{f}}.$ Following the same procedure as stated for the altitude, the true fuel level of the spacecraft is binned into the corresponding fuel band.
@@ 4.3 Reward Modeling
At a high level, the objective of the GRACE-FO mission is to provide high-quality observations of changes in Earth’s gravitational field [17]. The quality of these observations generally increases with lower atmospheric density, which occurs at higher altitudes. Due to this fact, the reward function $R$ should be independent of the choice of the action selection and only a function of the current state (i.e., altitude) and time step of the mission. The dependence of the reward function $R$ on time could be due to, e.g., time-varying atmospheric conditions that affect the observation quality. Thus, we introduce $R(s,:,h)\triangleq R(s,a,h)$ for all $a\in\mathcal{A}$ to denote the reward (a function of the observation quality) for being in a state $s\in\mathcal{S}$ corresponding to a fixed, discrete altitude. In other words, $R(s,:,h)$ is independent of the action taken. Note that $R(s,:,h)$ should be non-negative and monotonically increasing as the corresponding altitude of $s$ increases.
An additional hindrance to observation quality is the existence of resonant orbits, which are orbits with a periodic ground track [18]. Due to the periodicity of this ground track, less of Earth’s surface is observed, reducing the quality of the global gravitational field estimation. Thus, for any discrete altitude state $s\in\mathcal{S}_{a}$ for which a resonant orbit exists in its corresponding altitude band $s-\frac{R_{a}}{2},s+\frac{R_{a}}{2}]$, we can set $R(s,:,h)=0$. Alternatively, given the periodicity of the resonant orbit $i$, $res_{i}$ (i.e., the rate at which the ground track repeats), one can scale the reward function by a multiplicative factor $\alpha<1$ to account for variations in the reduction of observation quality. The reward function can then be defined as
@@
where $\omega_{e}$ is the Earth’s rotation rate and $\dot{\Omega}$ is the nodal precession rate. Substituting these into the resonance condition yields
@@ 4.4 Action and Fuel Cost Modeling
There are several choices to model the discrete set of actions $\mathcal{A}$ for the GRACE-FO mission. At their core, these actions represent the choice of how much altitude to gain over the subsequent discrete time interval. For example, if the temporal discretization is fine enough, then a single, continuous burn could be approximated by a sequence of smaller burns performed at each discrete time step. In this case, the set of actions is $\mathcal{A}=\{no\,burn,{}^{+}$$R_{a}\}$, where $R_{a}$ denotes the choice to increase the altitude by $R_{a}$. Since the bands are assumed to be evenly spaced in altitude, this action corresponds to transitioning to the state $s^{\prime}\in\mathcal{S}$ such that the altitude of $s^{\prime}$ is equal to the altitude of $s$ plus $R_{a}$. For models with lower temporal resolution, multiple actions corresponding to a finite set of altitude increases can be included in the set $\mathcal{A}$, i.e., $\mathcal{A}=\{no\,burn,{}^{+}$$R_{a},{}^{+}$$2R_{a},\ldots,{}^{+}$$N_{A}R_{a}\}$, where $N_{A}R_{a}$, $N_{A}\in\mathbb{Z}$, is the maximum one-step orbit increase. A visualization of such a set $\mathcal{A}$ is shown in Figure 5, where there exists a set of seven possible actions (no burn and six possible altitude increases). The selected action is in black. A third option for the choice of action is to use a geometric set to approximate the choice of increasingly large fuel burns. Specifically, let $\hat{\gamma}>1$ be a constant scaling factor and $N_{A}$ be the number of actions. Then, $\mathcal{A}$ corresponds to the altitude increase $\{no\,burn,{}^{+}\lceil\hat{\gamma}^{0}\rceil R_{a},{}^{+}\lceil\hat{\gamma}^{1}\rceil R_{a},\ldots,{}^{+}\lceil\hat{\gamma}^{N_{A}-1}\rceil R_{a}\}$.
Associated with each action is a fuel cost, while the decision-maker must ensure that the cumulative fuel cost throughout the mission does not exceed its fuel budget. Let us denote the fuel cost when taking action $a\in\mathcal{A}$ while in a state $s\in\mathcal{S}$ by $C(s,a)$ and the fuel budget by $C_{\mathrm{max}}$. Trivially, for the action corresponding to no burn, the fuel cost is zero. For the remaining actions, we present two options for modeling the fuel costs and the fuel budget:
@@
    Maximum Altitude Change. As previously discussed, each action is associated with an increase in the operational altitude of the GRACE-FO mission. The cost of an action is its corresponding altitude increase, i.e., $C(s,{}^{+}$$nR_{a})={}^{+}$$nR_{a}$ for all $n\in\{1,\ldots,N_{A}\}$. Similarly, the fuel budget is a maximum cumulative altitude increase over the mission horizon.
@@ 4.5 Transition Modeling
[…]
@@
where $R_{s}$ is the altitude corresponding to state $s\in\mathcal{S}$ and $^{+}R_{a}$ is the altitude gain corresponding to selecting action $a\in\mathcal{A}$. Note that (12c) is obtained from (12b) by explicitly including the conditioning of the transition probability on the solar flux. Furthermore, since the state dynamics themselves are assumed to have no stochastic disturbances, $\mathrm{Pr}(\mathrm{alt}_{t+1}\in[R_{s^{\prime}}-\frac{R_{a}}{2},R_{s^{\prime}}+\frac{R_{a}}{2}]|\mathrm{alt}_{t}=R_{s}+{}^{+}R_{a},f)$ takes either a value of 0 or $1$ (i.e., it is binary). In practice, the integral over $f$ can be numerically approximated, and the state transition matrix can be used to efficiently compute the subsequent altitude. A visualization of an example trajectory (with respect to the mission altitude) is shown in Figure 6.
@@
Note that decisions are only made at the discrete time instances represented by the densely dotted vertical lines.
@@
    Collision avoidance. When the risk of collision with another spacecraft or orbital debris increases, GRACE-FO cannot execute effective avoidance maneuvers. This risk is quantified by a collision rate, $\lambda_{c}$, whose integration over discrete time intervals yields the probability that a collision avoidance maneuver is required.
@@ 4.6 Problem Formulation
We now formalize the mission planning problem as an optimization task over a finite planning horizon. The objective is to maximize the total expected reward while ensuring that a critical safety specification is satisfied with high probability. We define a safety requirement expressed as a temporal logic formula that captures two key constraints:
@@
whereR(s,a,h)R(s,a,h) is the reward obtained at time step $h$, $s_{0}$ is the initial state, $\varphi$ is the temporal logic safety specification, and $\delta$ denotes the maximum allowable probability of a safety violation. The informal safety specification can be read as follows:
“The spacecraft does not perform a maneuver more often than every 3 months, and it never drops below 300 km.”
@@ 5.2 Planning Under Uncertainties
Next, we incorporate uncertainties into the model by treating both the solar flux and the outcomes of the orbit-raising maneuvers as random variables. Figure 10 shows the evolution of the solar flux over time, with the 75% quartile range highlighted in orange; these discretized distributions are integrated into the model. We then evaluate state transitions by considering best-case, worst-case, and average scenarios based on variations in solar flux and orbit-raising cost. Figure 10 displays a range of possible trajectories for altitude and fuel consumption. As expected, the worst-case scenario leads to fuel depletion early in 2025, while the best-case scenario extends fuel availability until 2029. The final altitude varies widely, from below 300 km up to 475 km, with an average of 410 km. Although the average behavior appears to satisfy the temporal specification, this plot alone does not provide full probabilistic safety guarantees. The results suggest possible violations of the requirement to maintain an altitude above 300 km, motivating us to turn next to safety verification.
@@
Figure 9: Altitude and fuel evolution when solving the MDP with uncertain transitions
@@
Figure 10: Probabilistic verification for the final altitude following the policy $\pi$ using the initial condition of GRACE-FO at its launch
@@ 5.3 Safety Verification
These results underscore the strength of the MDP framework in enforcing strict safety constraints while optimizing performance. Unlike traditional planning methods that do not explicitly verify safety through continuous feedback, the value function in this approach provides a clear measure of safety compliance. The near-zero risk of dropping below the minimum altitude and the high probability of maintaining an altitude above 425 km indicate that the policy not only meets but exceeds safety requirements.
@@
Figure 11: Altitude and fuel evolution when solving the MDP with uncertain transitions using GRACE-FO state in March 2025 as the initial condition
@@ 5.4 Current-State Planning
We can also use the previously synthesized policy to update the mission forecast by re-simulating from the spacecraft’s current state, reducing uncertainty about future trajectories. For instance, in early 2025, with only 1.5 kg of propellant remaining, propagating the existing policy yields a maneuver sequence in Fig. 11 that maintains altitude safely above 300 km, and in fact no scenarios fall below 400 km. This demonstrates how the method can incorporate real mission feedback to refine expectations about future performance under the existing policy.
In addition, we can resolve the MDP when improved solar-activity forecasts become available. By updating the transition probabilities, this would enable the resulting policies to adapt to evolving environmental predictions, reducing conservatism compared to fixed-scenario approaches and improving robustness under uncertainty.